Details.

Sky cloudy

Open chamber 5PM

OSCO preset for collimation.

opening 5:19PM

preset and not star in the guider.

Try to acquire a bright star.

10 magnitide, collimated.

osco

5:50pM

TN: NCPA OFF: 20211123_010059

NCPA ON.

TN:_010802

Stop thick clouds passing not data acauired.

osco_no_ncpa.png

6:24OPM too think...

magnitude 14....

magnitiude-20211122.png

6:43PM
  • cancel preset.
  • closing dome for Argos checkout.
  • There is a band of clouds passing.....

CHECK PUPIL bin4

Tue Nov 23 01:56:29.938 2021 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[24] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:-2.99182[-2.99182],soy:2.99182,rox:0[0],roy:0,mag:17,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Tue Nov 23 01:56:29.953 2021 aos.info - WFS source set to: FLAOWFS

rotation started 7:07PM

@180 cl x->70...

Moving forward with the pupil shifted.,..

loaded pupil #23

8:20PM

better all the rotation. 360 to 0

8:36PM back to open.

pupil20211117_192657.png

open chamber.

presetAO

Tue Nov 23 03:45:27.017 2021 aos.PresetAO{FLAO}.failed - PresetAO {FLAO}[29] - [AOException] presetAO: FATAL - Timeout waiting for techviewctrl.L.STATUS (-5001) TIMEOUT_ERROR
Tue Nov 23 03:46:28.480 2021 aos.Stop{FLAO}.started - Stop{FLAO}[30] (msg:Preset canceled)
Tue Nov 23 03:46:28.498 2021 aos.Stop{FLAO}.complete - Stop{FLAO}[30]

rotator in red->change rotator angle.

cloudy... preset not star found...

9:29 PM Argos cal unit

Tue Nov 23 04:32:03.955 2021 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[38] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:-2.99182[-2.99182],soy:2.99182,rox:0[0],roy:0,mag:17,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Tue Nov 23 04:32:03.969 2021 aos.info - W

guide probe park.

open loop

load shape.

the pupils loos bad.

open loop try bin1

TN:20211123_050639

psf_20211123_closedome.png

ummmm

rotation test to define the center...

limits: 176.85,181.35

optimize gain: 1.6 0.8,0.8

re-rotator position:179.1

optimize gain.

1.96,0.18,0.36

modify the wfsarb to intorduce the rotation offset.

20211123_052154

psf_20211123_b1-closeloop.png

preset AO for bin4

Tue Nov 23 05:42:21.637 2021 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[48] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:-2.99182[-2.99182],soy:2.99182,rox:0[0],roy:0,mag:17,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

same pdf..

homing ADC

Tue Nov 23 05:52:47.875 2021 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[51] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:-2.99182[-2.99182],soy:2.99182,rox:0[0],roy:0,mag:17,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

tn:_055754

same shape

TTM

Tue Nov 23 06:00:28.107 2021 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[53] (aomode:TTM-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:-2.99182[-2.99182],soy:2.99182,rox:0[0],roy:0,mag:17,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

tn:_060607

looks the same.

presetAO

Tue Nov 23 06:11:59.910 2021 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[55] (aomode:TTM-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:-2.99182[-2.99182],soy:2.99182,rox:0[0],roy:0,mag:17,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

Another

Check cube translator, rotator..

_062955

check hexapod correction

Tue Nov 23 06:32:39.114 2021 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[60] (aomode:TTM-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:-2.99182[-2.99182],soy:2.99182,rox:0[0],roy:0,mag:17,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

same issue

Tue Nov 23 06:42:03.333 2021 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[62] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:-2.99182[-2.99182],soy:2.99182,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

No disturbance

lngain.

tn:_064602

optimize gain:

Tn:_064849

psf_20211123_bin1_2.png

gopt:

_065435

gain zero

_065635

ace-bin 4

lngain

070616

manual gain:

tt0.2,0.15,0.05

tn:070944

luci dark:

11-22 23:45:28.001098| MAIN > arbFunc updateModulationInternal (, 'FLAOWFS', 'LUCIFER', 1000

presetb4

nModes:54 r0:nan snMode:0 strehl:0 ttMod:0
Tue Nov 23 07:26:44.216 2021 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[66] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:-2.99182[-2.99182],soy:2.99182,rox:0[0],roy:0,mag:17,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Tue Nov 23 07:26:44.232 2021 aos.info - WFS source set to: FLAOWFS

forgetting factorBi4soul

doen't show in changes.

TTM bin1.

_075218

10l/d

_080030

disturbance disbaled.

080418

bin2

Tue Nov 23 08:12:25.971 2021 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[72] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:-2.99182[-2.99182],soy:2.99182,rox:0[0],roy:0,mag:14,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

psf_20211123_bin2.png3| -13006 - [AOException] RETRY: Optical gain is not one, cannot start tracking (-5002) VALUE_OUT_OF_RANGE_ERROR

tn:_082157

ncpa on

_082604

psf_20211123_bin2_2.png

_084423

2:20am

warmup ocam.

power off/on

porcess stop/start

ace-b4

close loop

lngain:

_095123 blind/blind

_095337

optimize gain not working...

manual gain:

tn:_095747

SAVED A FLAT.

  • 20211123.SAV
BIN 1

PRESETAO BIN4

Tue Nov 23 10:35:12.849 2021 aos.PresetAO{FLAO}.started - PresetAO{FLAO}[85] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:-2.99182[-2.99182],soy:2.99182,rox:0[0],roy:0,mag:17,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Tue Nov 23 10:35:12.863 2021 aos.in

PSF:

20211123_104258

close loop b4

LNgain.

_105210

move rerotator.

182.25 bad...

174.25 bad
  • _105908
move to 178.25

wfsarb stop/start

turn on rerotator->179.1
  • tn:_110348
optimize gain.

failed.

loaded dist file.

it didn't work.

Rotator offset.

RotatorOffsetBin1 float 349.6 # Should 350.75 to match with REC 20181215_201757. DLM Adjusted 21211122 after changing homing speed

# last used value:
# RotatorOffsetBin1 float 215.25 # Setting for telescope derotator=0, Cold coolant temps (1.7C)
#RotatorOffsetBin1 float 226.2 # Setting for telescope derotator=0, Cold coolant temps (-7.7C)
RotatorOffsetBin2 float 350.25
RotatorOffsetBin3 float 350.4
#RotatorOffsetBin4 float 225.4
#RotatorOffsetBin4 float 350.25
RotatorOffsetBin4 float 348.75 #DLM Adjusted 21211122 after changing homing speed

STOP.LOADED THE PUPIL BUILT WITH THE RECONSTRUCTOR.

move to 270

TTM table link to .20211025 where the Bin4 pupils were 20200916_173857

rwxrwxrwx. 1 AOeng AOeng 33 Nov 23 11:28 table_LUCIFER_TTM-AO.txt -> table_LUCIFER_TTM-AO.txt.20211025
lrwxrwxrwx. 1 AOeng AOeng 33 Nov 23 11:38 table_LUCIFER_ACE-AO.txt -> table_LUCIFER_ACE-AO.txt.20211117

change split to x=1,y=0
  • TN: 20211123_120547
  • looks ground

back to today pupils.

split x=3, y=1.

pupil diam 10.1

stop....

4:45am

WFS left power ON.

-- JuanCarlosGuerra - 23 Nov 2021
Topic revision: r3 - 23 Nov 2021, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback